When you retire you find that you have $1,000,000 in your retirement account into which you have been depositing $750 each month for the last 40 years. What annual rate did your money earn? Round your answer to 2 decimals.
Answer #1

Using financial calculator
Input:

FV = 1000000

PMT= -750

N = 40*12= 480

Solve for I/Y as 0.3731

Annual rate = 0.3731*12 = 4.48%